# NON-FACTORIZABLE RIBBON HOPF ALGEBRAS

### v1

@@ 5.3 Cases of degeneration of the invariant
When the unimodular ribbon Hopf algebra is factorizable Beliakova and De Renzi proved that the invariant we obtained is actually an invariant of the boundary of the handlebody. We call this situation degeneration to dimension $3$.
@@ Appendix A Proofs
Here we used the fact the the map $i$ of Remark 4.10 is an algebra map which can be checked directly. The list of the relations between the generators is an exhaustive one, as the choices of action and coaction force them to be diagonal. Moreover, with the operations defined for arbitrary monomials as above, we see that the new relations respect the monomial bases of both pieces. Since both $U$ and $H$ are fully defined by the relations on generators and their extension over the PBW- and monomial bases respectively, and this fact is respected by the new operations, we can conclude there are no relations between the pieces $U$ and $H$ not resulting from the relation of Equation 3. That is, we obtain the exhaustive list of relations of $U\ltimes H$ by deriving them from the relations between the generators and applying the fact they respect the new operations of the Hopf algebra.
@@
$=$ $\Delta(E^{a}F^{b}K^{c}E^{e}F^{f}K^{d}\otimes\mathbf{K}^{\mathbf{v}}\mathbf{X}^{\mathbf{r}})$
@@
$=$ $\Delta(E^{a}F^{b}K^{c}\otimes\mathbf{K}^{\mathbf{v}_{1}}\mathbf{X}^{\mathbf{r}_{1}}\mathbf{K}^{\mathbf{v}_{2}}\mathbf{X}^{\mathbf{r}_{2}})$
@@
It is straightforward to find the Drinfeld element of $R=DR_{\boldsymbol{\mathsf{z}}}\Theta\bar{R}_{\boldsymbol{\alpha}}$ to be $u:=u_{U}u_{H}$ since the two pieces coming from the respective algebras commute and the piece coming from $\bar{R}_{\boldsymbol{\alpha}}$ is in fact just $u_{H}$ as for $R_{\boldsymbol{\alpha}}$, because $(K^{r^{\prime\prime}})^{2}=\mathbf{1}$. We need to find a pivotal element $g\in G(U\ltimes H)$ that would make the ribbon element $v:=g^{-1}u$ central. This is achieved by choosing $g=K$, as it verifies the criterion for $K,E,F$ by assumption, and $\{K,X_{k}\}=0$, which is the criterion necessary for $S^{2}(X_{k})=KX_{k}K^{-1}=-X_{k}$. The ribbon axioms can be checked directly, using the relations in $U\ltimes H$ and the fact that both $U,H$ carry ribbon structures that we know explicitly. Let us denote them $v_{H}$ and $v_{U}$. Since $S(v_{H})=v_{H}$ and $S(g_{H})=g_{H}$, we have $S(u_{H})=u_{H}$, because $[g_{H},u_{H}]=0$. Then we have
@@ B.2 Non-factorizable extensions of $u_{q}\mathfrak{sl}_{2}$ examples
$\Delta(K_{1})=K_{1}\otimes K_{1},\;\;\Delta(Z^{\pm})=\mathbf{1}\otimes Z^{\pm}+Z^{\pm}\otimes L^{\pm 1}.$

### v2

@@ 5.3 Cases of degeneration of the invariant
When the unimodular ribbon Hopf algebra is factorizable Beliakova and De Renzi proved that the invariant we obtained is actually an invariant of the boundary of the handlebody together with its signature. We call this situation degeneration to dimension $3$.
@@ Appendix A Proofs
Here we used the fact the the map $\iota$ of Remark 4.10 is an algebra map which can be checked directly. The list of the relations between the generators is an exhaustive one, as the choices of action and coaction force them to be diagonal. Moreover, with the operations defined for arbitrary monomials as above, we see that the new relations respect the monomial bases of both pieces. Since both $U$ and $H$ are fully defined by the relations on generators and their extension over the PBW- and monomial bases respectively, and this fact is respected by the new operations, we can conclude there are no relations between the pieces $U$ and $H$ not resulting from the relation of Equation 3. That is, we obtain the exhaustive list of relations of $U\ltimes H$ by deriving them from the relations between the generators and applying the fact they respect the new operations of the Hopf algebra.
@@
$=$ $\Delta(E^{a}F^{b}K^{c}E^{e}F^{f}K^{d}\otimes\mathbf{K}^{\mathbf{v}}\mathbf{X}^{\mathbf{r}}).$
@@
$=$ $\Delta(E^{a}F^{b}K^{c}\otimes\mathbf{K}^{\mathbf{v}_{1}}\mathbf{X}^{\mathbf{r}_{1}}\mathbf{K}^{\mathbf{v}_{2}}\mathbf{X}^{\mathbf{r}_{2}}).$
@@
It is straightforward to find the Drinfeld element of $R=DR_{\boldsymbol{\mathsf{z}}}\Theta\bar{R}_{\boldsymbol{\alpha}}$ to be $u:=u_{U}u_{H}$ since the two pieces coming from the respective algebras commute and the piece coming from $\bar{R}_{\boldsymbol{\alpha}}$ is in fact just $u_{H}$ as for $R_{\boldsymbol{\alpha}}$, because $(K^{r^{\prime\prime}})^{2}=\mathbf{1}$. We need to find a pivotal element $g\in G(U\ltimes H)$ that would make the ribbon element $v:=g^{-1}u$ central. This is achieved by choosing $g=K$, as it verifies the criterion for $K,E,F$ by assumption, and $\{K,X_{k}\}=0$, which is the criterion necessary for $S^{2}(X_{k})=KX_{k}K^{-1}=-X_{k}$. The ribbon axioms can be checked directly, using the relations in $U\ltimes H$ and the fact that both $U,H$ carry ribbon elements that we know explicitly. Let us denote them $v_{H}$ and $v_{U}$. Since $S(v_{H})=v_{H}$ and $S(g_{H})=g_{H}$, we have $S(u_{H})=u_{H}$, because $[g_{H},u_{H}]=0$. Then we have
@@ B.2 Non-factorizable extensions of $u_{q}\mathfrak{sl}_{2}$ examples
$\Delta(K_{1})=K_{1}\otimes K_{1},\;\;\Delta(Z^{\pm})=\mathbf{1}\otimes Z^{\pm}+Z^{\pm}\otimes L^{\pm 1}$
@@
$\Delta(K_{1})=K_{1}\otimes K_{1},\;\;\Delta(K_{2})=K_{2}\otimes K_{2},\;\;\Delta(X^{\pm})=\mathbf{1}\otimes X^{\pm}+X^{\pm}\otimes L^{\pm 1}$